Pitchfork, Inc. is preparing its 2020 financial statements. The company's accountant calculated Income from Continuing Operations to be $1,700,000, but upon further review is not certain this number is accurate. Pitchfork has a corporate income tax rate of 30%. Additionally, the company reports only one year of financial data on the face of the financial statements.  All amounts listed are pretax unless otherwise noted. After reviewing the following information, determine the appropriate adjustments, if any, to Income from Continuing Operations. Once you have determined the CORRECT Income from Continuing Operations, complete the remainder of the Income Statement for reporting EPS.

1. On January 1, 2017, Pitchfork purchased a machine for $180,000 with a salvage value of $20,000 and useful life of eight years which was depreciated using the straight-line method. During 2020, Pitchfork decided to change to double-declining-balance method. The $1,700,000 Income from Continuing Operations had already been calculated using the straight-line depreciation method.

Determine the correct ADJUSTMENT to Income from Continuing Operations (ICO) for Depreciation Expense in 2020.

Adjustment for Depreciation Expense (2020):___________

Continuing with the information presented in #1 above, Pitchfork has ICO of $1,700,000 and a corporate tax rate of 30%. Determine if ICO should be adjusted based on the following information:

2. Pitchfork had an unrealized loss from foreign currency translation adjustments of $120,000 (pretax) that was included in calculating the $1,700,000 income from continuing operations.

Adjustment to I.C.O. for Translation Loss from Foreign Currency: __________

Continuing with the information presented in #1 above, Pitchfork Inc has Income from Continuing Operations (ICO) of $1,700,000 and a corporate tax rate of 30%. Determine if ICO should be adjusted based on the following information:

3. During 2020, Pitchfork closed one of its stores for a pre-tax loss of $150,000. This store closure did not qualify as a component of the entity, nor did it create a strategic shift in the operations of the entity. Therefore, it should not be treated as Discontinued Operations. The $150,000 restructuring charges were excluded in determining the $1,700,000 income from continuing operations.

To correct I.C.O., the Adjustment for Restructuring Charges would be $ _________

Continuing with the information presented in #1 above, Pitchfork has Income from Continuing Operations (ICO) of $1,700,000 and a corporate tax rate of 30%. Determine if ICO should be adjusted based on the following information:

4. On April 1, 2019 Pitchfork paid $24,000 for two years rent on office space and at the time debited Rent Expense. No adjusting or correcting entries were made for this transaction in 2019 or 2020.

a. To correct I.C.O for 2020, the correct Rent Expense (after tax) would be: $ _________


b. Determine the amount of the Prior Period Adjustment to be reported on the Retained Earnings Statement to correct the Beginning Balance at Jan 1, 2020: ______

5. Pitchfork sold investments during the year that resulted in a pre-tax loss of $18,000. The company also had unrealized gains on Available for Sale securities of $20,000 (pre-tax). Both of these transactions were excluded in determining the $1,700,000 Income from Continuing Operations calculation.

To correct I.C.O. for 2020, the adjustment for gains/losses on investments would be:   $_________

6. Using the adjustments you made in items 1-5 above, determine the CORRECTED Income From Continuing Operations _________

7. Referring to the information presented above in questions 1-6, determine Pitchfork's Comprehensive Income as of year-end: $_________
Answer #1

1. Adjustment for Depreciation Expense (2020):$ (19,600)

Straight line depreciation for the first three years = {$ ( 180,000 - 20,000 ) / 8 } * 3 = $ 60,000.

Book value as on Jan 1, 2020 = $ 180,000 - $ 60,000 = $ 120,000.

Double declining rate of depreciation = 100 / 5 * 2 = 40 %.

Double declining depreciation expense for 2020 = $ 120,000 x 40 % = $ 48,000.

Depreciation already provided = $ 20,000.

Further depreciation expenses needed to be recognized = $ 28,000

Adjustment needed = $ 28,000 ( 1 - 0.3) = $ 19,600.

2.Adjustment to I.C.O. for Translation Loss from Foreign Currency: $ 84,000

Unrealized loss on foreign currency translations are to be reported on the Statement of Other Comprehensive Income. Therefore adjustment required = $ 120,000 * ( 1 - 0.30) = $ 84,000.

3. To correct I.C.O., the Adjustment for Restructuring Charges would be $ ( 105,000)

Adjustment needed = $ 150,000 * ( 1 - 0.3 ) = $ 105,000

4.a. To correct I.C.O for 2020, the correct Rent Expense (after tax) would be: $ ( 8,400)

Rent expense for 2020 = $ 12,000

Adjustment needed = $ 12,000 * ( 1 - 0.30) = $ 8,400.

b. Determine the amount of the Prior Period Adjustment to be reported on the Retained Earnings Statement to correct the Beginning Balance at Jan 1, 2020: $ 10,500

Unexpired rent as of Jan 1, 2020 = $ 24,000 - $ 9,000 = $ 15,000.

Therefore adjustment needed to Retained Earnings = $ 15,000 * ( 1 - 0.3) = $ 10,500.

5.

To correct I.C.O. for 2020, the adjustment for gains/losses on investments would be: $ (12,600)

Adjusted pre-tax realized loss on investments = $ 18,000 * ( 1 - 0.3) = $ 12,600.

Unrealized gain on available for sale securities has been correctly excluded.